john caught 60 fish on saturday and tagged them all, on tuesday he caught 60 fish, three were tagged. he returned all to the pond. if the percent of tagged fish that john caught on tuesday equals the percent of all the tagged fish in the pond, what is the number of fish in the pond
Answer by stanbon(75887)

john caught 60 fish on saturday and tagged them all, on tuesday he caught 60 fish, three were tagged. he returned all to the pond. if the percent of tagged fish that john caught on tuesday equals the percent of all the tagged fish in the pond, what is the number of fish in the pond
---------
Let "x" be the # of fish in the pond.
--------
x/60 = 60/3
x = 60^2/3 = 20*60 = 1200
======================
Cheers,
Stan H.
======================